OpenStudy (eric_d):

solve following equation for values of theta in range of................., in terms of pi

OpenStudy (eric_d):

 @ikram002p

OpenStudy (eric_d):

\[7\tan2x = 4sinx\]

OpenStudy (eric_d):

Try solving this equation manually

OpenStudy (eric_d):

The book answer is 0, 0.580pi, pi, 1.420pi, 2pi

OpenStudy (eric_d):

@PFEH.1999

OpenStudy (ikram002p):

is it like this ? \(\Huge 7 \tan 2x = 4 \sin x \) for 0≤x≤2π

OpenStudy (eric_d):

yes

OpenStudy (ikram002p):

nw tan 2x= ?

OpenStudy (ikram002p):

it has identity do you know it ?

OpenStudy (ikram002p):

\(tan(2x) = [2 tan(x)] / [1 - tan^2(x)] \)

OpenStudy (eric_d):

ok..

OpenStudy (ikram002p):

tan(2x) = [] / [1 - tan^2(x)] \[\Huge tan(2x) =\dfrac { 2 \tan(x) } { 1-\tan^2 (x) }\]

OpenStudy (ikram002p):

so \(\Huge 4\sin x =7 \dfrac { 2 \tan(x) } { 1-\tan^2 (x) }\)

OpenStudy (ikram002p):

\(\Huge 4\sin x(1-\tan^2 (x) ) = 14 \tan(x) \)

OpenStudy (eric_d):

ok..

OpenStudy (ikram002p):